OpenStudy (revircs):

Need guidance in differentiating f(x)=arctan(√(ln(1/x))

OpenStudy (revircs):

\[\arctan(\sqrt{\ln(\frac{ 1 }{ x }})\]

OpenStudy (astrophysics):

\[\frac{ d }{ dx } (\tan^{-1}x)=\frac{ 1 }{ 1+x^2 }\] so you will have to use chain rule here, is your derivative \[ \frac{ d }{ dx } \arctan \left( \sqrt{\ln (\frac{ 1 }{ x })}\right)\]

OpenStudy (astrophysics):

I think it might be easier if you like \[\sqrt{\ln \frac{ 1 }{ x }} = u\] and then make the substitution back so you can see what exactly is going on

OpenStudy (astrophysics):

let*

OpenStudy (astrophysics):

Or you can just take the derivative haha, it's going to be ugly, try using the chain rule and see what you get

OpenStudy (revircs):

so I would get \[\frac{ 1 }{ \ln \frac{ 1 }{ x } } x \frac{ d }{ dx }\sqrt{\ln \frac{ 1 }{ x }}\], then the chain rule again?

OpenStudy (revircs):

sorry, 1/(1+ln(1/x))

OpenStudy (astrophysics):

\[\frac{ d }{ dx } \arctan \left( \sqrt{\ln (\frac{ 1 }{ x })}\right) = \frac{ 1 }{ 1+\ln(1/x) } * \ln(\frac{ 1 }{ x })'\] do you mean this

OpenStudy (revircs):

Yeah. Im not really fimiliar with the tools yet, haha

OpenStudy (revircs):

Then since its the derivative of ln(1/x), we would have to apply the chain rule again?

OpenStudy (astrophysics):

One sec it's not as easy as that, let me try on paper and I'll get back to you in a second

OpenStudy (astrophysics):

Ok so we should get \[- \frac{ 1 }{ \sqrt{\ln \left( \frac{ 1 }{ x } \right)}(2x+2xln \left( \frac{ 1 }{ x } \right) }\] it's probably easier making the u substitution and breaking it in parts, or we can do it all at once which is kind of nasty

OpenStudy (astrophysics):

\[\frac{ d }{ dx } \left( \sqrt{\ln(\frac{ 1 }{ x })} \right)\] so what's the derivative of this?

OpenStudy (revircs):

i believe it would be \[\frac{ -1 }{ 2x \sqrt{\ln \frac{ 1 }{ x }} }\]

OpenStudy (astrophysics):

Right, awesome!

OpenStudy (astrophysics):

Ok now we have \[\frac{ d }{ dx } arctanu= \frac{ 1 }{ 1+u^2 }\] so now we just take \[u = \sqrt{\ln(1/x)}\] and substitute that in, and we can now multiply by that derivative above and we're done :)

OpenStudy (astrophysics):

Does that make sense?

OpenStudy (revircs):

kind of, do we not need to use the chain rule again for 1/x?

OpenStudy (astrophysics):

Only reason I broke it up is because we would be hesitant in taking the derivative of just ln(1/x) which is wrong

OpenStudy (astrophysics):

We did take the derivative, we took the derivative of \[\sqrt{\ln(1/x)}\] which included that

OpenStudy (astrophysics):

Ok I guess I'll show the long ugly way to, \[\frac{ d }{ dx } \arctan(\sqrt{\ln(1/x)}) = \frac{ 1 }{ 1+\left( \sqrt{\ln(1/x)} \right)^2 } \times [\sqrt{\ln(1/x)}]'\] I guess it's not too bad but you can't square for the derivative

OpenStudy (revircs):

OH, duh so we get down to \[\frac{ 1 }{ 1+\ln \frac{ 1 }{ x } }*\frac{ -1 }{ 2x \sqrt{\ln \frac{ 1 }{ x }} }\]

OpenStudy (astrophysics):

Yes, exactly :) so we then have \[-\frac{ 1 }{ 2xln \sqrt{1/x}(1+\ln(1/x)) }\]

OpenStudy (revircs):

Is this the same as the one you had above?

OpenStudy (astrophysics):

Yup, it doesn't matter

OpenStudy (revircs):

Oh, okay awesome. Thank you for your help!

OpenStudy (astrophysics):

The above just multiplied through by 2x